a simple random sample of 26 college students showed a mean credit score of 575 and a standard deviation 10. construct the 90% confidence interval estimate of the mean credit score for all college students

a. state the critical value

b. compute the margin of error

c. state the confidence interval
Answer #1

SOLUTION:

From given data,

a simple random sample of 26 college students showed a mean credit score of 575 and a standard deviation 10. construct the 90% confidence interval estimate of the mean credit score for all college students.

\bar{x} = 575

s = 10

n = 26

Degree of freedom = df = n-1 = 26-1 = 25

a. state the critical value

90% confidence interval

c = 90% = 90/100 = 0.9

\alpha = 1-c = 1-0.9 = 0.1

\alpha/2 = 0.1/2 = 0.05

critical value

t\alpha/2,df = t0.05,25 = 1.708

b. compute the margin of error

Margin of error = E = t\alpha/2,df *(s/sqrt(n))

E =1.708*(10/sqrt(26))

E =1.708*1.961161

E = 3.3497

c. state the confidence interval

\bar{x} - E < \mu < \bar{x} + E

575 - 3.3497 < \mu < 575 + 3.3497

571.6503 < \mu < 578.3497

571.65 < \mu < 578.35

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.